# appendixReferences

### v3

@@ A.3 Proof of Auxilliary Lemmas
$\leq 1+\mathbb{E}\left[\frac{|\bar{S}|-\hat{N}_{\bar{S}}}{N}|F_{S}\right]+\mathbb{E}\left[\frac{N_{Q}-\hat{N}_{Q}}{N}|F_{S}\right]+\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S}\right]$
$\leq 1+\frac{|\bar{S}|}{Np_{\min}}+0+\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=0\right]+P(B=1|F_{S})\cdot\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=1\right]$ (115)
$\leq 1+\frac{|\bar{S}|}{Np_{\min}}+\frac{(|\bar{S}|+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})p_{\min}}+\frac{1}{N^{3}}\cdot\left(\frac{N}{p_{\min}}\right)^{2}\cdot\frac{1}{N}$ (116)
$=O\left(\frac{1}{N}\right)+O\left(\frac{\log N}{N}\right)+O\left(\frac{1}{N^{2}}\right)$ (117)
where (114) holds by the identity $\frac{a}{b}=1+\frac{a-b}{a}+\frac{(a-b)^{2}}{ab}$; (115) holds because conditioned on $F_{S}$, it holds that $1\leq\hat{N}_{\bar{S}}\leq\frac{|\bar{S}|}{p_{\min}}$; (116) holds by definition of the event $\{B=0\}$ and because conditioned on $F_{S}$, it holds that $1\leq\hat{N}_{Y_{ki}}\leq N/p_{\min}$; and (117) holds because $|\bar{S}|\leq 2d_{\max}^{2}$, $\epsilon_{B}=\sqrt{3C^{2}N\log N}$, $N_{Q}\geq N-2d_{\max}$, and (113). Letting $S=\{i\}$ implies (62), while letting $S=\{i,j\}$ implies (63).
@@
$\leq\left(\frac{|\bar{S}|}{Np_{\min}}\right)^{2}+\frac{Nd_{\max}}{p_{min}^{2}N^{2}}+\mathbb{E}\left[{\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right)|F_{S}\right]+$
$\hskip 28.45274pt{\operatorname{Var}\,}\left(\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right]|F_{S}\right)$
$\leq\left(\frac{|\bar{S}|}{Np_{\min}}\right)^{2}+\frac{Nd_{\max}}{p_{min}^{2}N^{2}}+\left(\frac{(|\bar{S}|+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}\right)^{2}+\left(\frac{N^{2}}{p_{\min}^{2}N}\right)^{2}\cdot\frac{1}{N^{3}}$
$=O\left(\frac{1}{N^{2}}\right)+O\left(\frac{1}{N}\right)+O\left(\frac{\log N}{N}\right)+O\left(\frac{1}{N}\right)$
@@
$\leq\sum_{i=1}^{3}{\operatorname{Var}\,}a_{i}+\sum_{i=1}^{3}\sum_{j:j\neq i}({\operatorname{Var}\,}a_{i}+{\operatorname{Var}\,}a_{j})$
and the remaining steps follow by similar reasoning as (115)-(117). ∎
@@
Let $h_{i}$ be defined as in Lemma 9, and let $\sigma^{2}={\operatorname{Var}\,}\sum_{i}h_{i}$. Divide the sequence of experiments into two subsequences, depending on whether $\sigma^{2}\geq N^{2/3+2\epsilon}$.
For the subsequence that satisfies $\sigma^{2}\geq N^{2/3+2\epsilon}$, it holds by linearization arguments made in the proof of Theorem 4 that
$\frac{\hat{\Phi}_{k}^{{\textup{Haj}}}-\Phi}{\sigma}$ $=\frac{\nabla h(\mathbb{E}\xi)^{T}(\xi-\mathbb{E}\xi)}{\sigma}+O_{P}\left(\frac{1}{\sigma}\right)$ (119)
@@
$\frac{\hat{\Phi}_{k}^{{\textup{Haj}}}-\Phi}{\sigma}\rightarrow N(0,1)$
@@
$\liminf\limits_{N\rightarrow\infty}{\operatorname{Var}\,}\left[\frac{\hat{\Phi}_{k}^{\textup{Haj}}-\Phi}{\sigma}\right]\geq 1$ (121)
@@
$\frac{V_{k}}{\sigma^{2}}$ $=\frac{\mathbb{E}V_{k}}{\sigma^{2}}+O_{P}\left(\frac{N^{1/2+\delta}}{\sigma^{2}}\right)$
$\geq{\operatorname{Var}\,}\left(\frac{\hat{\Phi}_{k}^{\textup{Haj}}}{\sigma}\right)+O_{P}\left(\frac{N^{1/2+\delta}}{\sigma^{2}}\right)$ (122)
@@
$\liminf\limits_{N\rightarrow\infty}\frac{V_{k}}{\sigma^{2}}\geq 1$
@@
$P\left(\Phi-\hat{\Phi}_{k}\geq z_{1-\frac{\alpha}{2}}\sqrt{V_{k}}\right)\leq\frac{\alpha}{2}+o(1)$ (123)
for the subsequence of experiments that satisfies $\sigma^{2}\geq N^{2/3+2\epsilon}$.
@@
$P\left(\left|\sum_{i}h_{i}\right|\geq\frac{N^{1/3+\epsilon}}{\sqrt{\alpha/2}}\right)\leq\alpha/2$
where we have used that $\sigma^{2}<N^{2/3+2\epsilon}$. In turn, this implies for any $\epsilon^{\prime}>0$ that that
$P\left(\hat{\Phi}_{k}-\Phi\geq(1+\epsilon^{\prime})\frac{N^{1/3+\epsilon}}{\sqrt{\alpha/2}}\right)\leq 1-\alpha+o(1)$ (124)
where we have used that $\hat{\Phi}_{k}-\Phi=\sum_{i}h_{i}+O_{P}(1)$, which holds by (119). In turn, we may choose $\epsilon^{\prime}$ and $\epsilon$ such that (124) is equivalent to
$P\left(\hat{\Phi}_{k}-\Phi\geq z_{1-\frac{\alpha}{2}}\sqrt{\Delta}\right)\leq\alpha/2+o(1)$ (125)
where $\Delta=(1.01\ N^{2/3+0.01})/(z_{1-\frac{\alpha}{2}}^{2}\cdot\alpha/2)$. Combining (123) and (125) yields that for all experiments, it holds that
$P\left(\hat{\Phi}_{k}-\Phi\geq z_{1-\frac{\alpha}{2}}\sqrt{\max\left(V_{k},\Delta\right)}\right)\leq\alpha/2+o(1),$
As $\widetilde{V}_{k}$ can be seen to equal $\max\left(V_{k},\Delta\right)$, this proves the lemma.
@@
It can be seen that $|h_{i}|\leq\frac{1}{p_{\min}}$, and that the dependency neighborhoods of $h_{1},\ldots,h_{N}$ have maximum size bounded by $d_{\max}^{2}$, since each $h_{i}$ depends on $X$ only through $X_{\Gamma_{i}}$. As a result, applying Theorem 7 yields
$d_{W}\left(\frac{\sum_{i}h_{i}}{\sigma}\ ,\ N(0,1)\right)$ $\leq\frac{d_{\max}^{6}N}{\sigma^{3}p_{\min}}+\frac{\sqrt{28}\,d_{\max}^{3}\,\sqrt{N}}{\sqrt{\pi}p_{\min}^{2}\sigma^{2}}$
@@
where (126) uses $\sigma^{2}\geq N^{2/3+\delta}$ for any $\delta>0$. ∎

### v4

@@ D.4 Proof of Theorem 7
$\leq 1+\left|\mathbb{E}\left[\frac{|\bar{S}|-\hat{N}_{\bar{S}}}{N}|F_{S}\right]\right|+\left|\mathbb{E}\left[\frac{N_{Q}-\hat{N}_{Q}}{N}|F_{S}\right]\right|+$
$\hskip 28.45274pt\left|\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S}\right]\right|$ (158)
$\leq 1+\frac{|\bar{S}|}{N\rho}+0+\left|\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=0\right]\right|+$
$\hskip 28.45274ptP(B=1|F_{S})\cdot\left|\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=1\right]\right|$ (159)
$\leq 1+\frac{|\bar{S}|}{N\rho}+\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}+\frac{1}{N^{2+\delta}}\cdot\left(\frac{N}{\rho}\right)^{2}\cdot\frac{1}{N}$ (160)
$=1+O\left(\frac{1}{M}\right)+O\left(\frac{\frac{N^{2}}{M^{2}}+\frac{N^{2}}{M}\log M}{N^{2}}\right)+O\left(\frac{1}{N^{1+\delta}}\right)$ (161)
$=1+O\left(\frac{\log M}{M}\right)$ (162)
where (157) holds by the identity $\frac{a}{b}=1+\frac{a-b}{a}+\frac{(a-b)^{2}}{ab}$; (158) decomposes $N-\hat{N}_{Y_{ki}}=|\bar{S}|-\hat{N}_{\bar{S}}+N_{Q}-\hat{N}_{Q}$; (159) uses the bound
$\left|\mathbb{E}\left[\frac{|\bar{S}|-\hat{N}_{|\bar{S}|}}{N}|F_{S}\right]\right|\leq\frac{|\bar{S}|}{N\rho},$
which holds because $0\leq\hat{N}_{\bar{S}}\leq\frac{|\bar{S}|}{\rho}$, and uses
$\mathbb{E}\left[\frac{N_{Q}-\hat{N}_{Q}}{N}|F_{S}\right]=0,$
which holds because $\mathbb{E}[\hat{N}_{Q}|F_{S}]=N_{Q}$; (160) uses $P(B=1|F_{S})\leq 1/N^{2+\delta}$ which holds by (156), and also uses
$\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}\leq\begin{cases}\displaystyle{\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}}&\textup{ if }B=0\\ \\ \displaystyle{\left(\frac{N}{\rho}\right)^{2}\cdot\frac{1}{N}}&\textup{ if }B=1\end{cases}$ (163)
which will be proven after the main result; (161) holds because $|\bar{S}|\leq\frac{2CN}{M}$ since $|S|\leq 2$, and also by definition of $\epsilon_{B}$, and $N_{Q}\geq N-\frac{2CN}{M}$. Letting $S=\{i\}$ implies (134), while letting $S=\{i,j\}$ implies (135).
@@
$\leq\left(\frac{|\bar{S}|}{N\rho}\right)^{2}+\frac{Cd}{\rho^{2}M}+\mathbb{E}\left[{\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right)|F_{S}\right]+$
$\hskip 28.45274pt{\operatorname{Var}\,}\left(\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right]|F_{S}\right)$ (165)
$\leq\left(\frac{|\bar{S}|}{N\rho}\right)^{2}+\frac{Cd}{\rho^{2}M}+\left(\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}\right)^{2}+\left(\frac{N^{2}}{\rho^{2}N}\right)^{2}\cdot\frac{1}{N^{2+\delta}}$ (166)
$=O\left(\frac{1}{M^{2}}\right)+O\left(\frac{1}{M}\right)+O\left(\left[\frac{\log M}{M}\right]^{2}\right)+O\left(\frac{1}{M}\right)$ (167)
$=O\left(\frac{1}{M}\right)$ (168)
@@
$\leq\sum_{i=1}^{3}{\operatorname{Var}\,}a_{i}+\sum_{i=1}^{3}\sum_{j:j\neq i}({\operatorname{Var}\,}a_{i}+{\operatorname{Var}\,}a_{j}),$
(165) uses the bounds
${\operatorname{Var}\,}\left[\frac{|\bar{S}|-\hat{N}_{\bar{S}}}{N}|F_{S}\right]$ $\leq\left(\frac{|\bar{S}|}{N\rho}\right)^{2}$ (169)
${\operatorname{Var}\,}\left[\frac{N_{Q}-\hat{N}_{Q}}{N}|F_{S}\right]$ $\leq\frac{CN^{2}d^{2}}{\rho^{2}MN^{2}}=\frac{Cd^{2}}{\rho^{2}M},$ (170)
which will be proven after the main result, and also uses the law of total variance to decompose ${\operatorname{Var}\,}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}\right]$; (166) uses the bounds
$\mathbb{E}\left[{\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right)|F_{S}\right]$ $\leq\left(\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}\right)^{2}+\frac{1}{N^{2+\delta}}\left(\frac{N^{2}}{\rho N}\right)^{2}$ (171)
${\operatorname{Var}\,}\left(\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right]|F_{S}\right)$ $\leq\frac{1}{N^{2+\delta}}\left(\frac{N^{2}}{\rho N}\right)^{2}$ (172)
which will be proven shortly; (167) holds because $|\bar{S}|\leq\frac{2CM}{N}$, by definition of $\epsilon_{B}$, and $N_{Q}\geq N-\frac{2CM}{N}$.
It remains to show (163), (169) – (170), and (171) – (172):
  1. 1.
    To show (163), we let $\textup{card}(\bar{S})$ denote $|\bar{S}|$, and observe that if $B=0$ then it holds that
    $\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}$ $=\frac{(|\bar{S}|-\hat{N}_{\bar{S}}+N_{Q}-\hat{N}_{Q})^{2}}{N(\hat{N}_{\bar{S}}+\hat{N}_{Q})}$
    $\leq\frac{(|\textup{card}(\bar{S})-\hat{N}_{\bar{S}}|+|N_{Q}-\hat{N}_{Q}|)^{2}}{N(\hat{N}_{\bar{S}}+\hat{N}_{Q})}$
    $\leq\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})},$ (173)
    where (173) uses that $|\hat{N}_{Q}-N_{Q}|<\epsilon_{B}$ when $B=0$, and also that $\hat{N}_{\bar{S}}\geq 0$ to lower bound the denominator, and also that $\hat{N}_{\bar{S}}\leq|\bar{S}|/\rho$ which implies $|\textup{card}(\bar{S})-\hat{N}_{\bar{S}}|\leq|\bar{S}|/\rho$.
    On the other hand, if $B=1$, it holds conditional on $F_{S}$
    $\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}$ $\leq\left(\frac{N}{\rho}\right)^{2}\cdot\frac{1}{N}$ (174)
    where the numerator uses $\hat{N}_{Y_{ki}}\leq N/\rho$ and the denominator uses that $\hat{N}_{Y_{ki}}\geq 1$ conditional on $F_{S}$. Combining (173) and (174) proves (163).
  2. 2.
    To show (169) we observe that $0\leq\hat{N}_{\bar{S}}\leq\frac{|\bar{S}|}{\rho}$.
  3. 3.
    To show (170) we observe that $\hat{N}_{Q}/N$ given by (154) is a sum of $\leq N$ terms whoses magnitudes are bounded by $(N\rho)^{-1}$ and whose dependency graph has at most $N\cdot\frac{d^{2}CN}{M}$ edges, where we have used (137).
  4. 4.
    To show (171), we observe that
    $\mathbb{E}\left[{\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right)|F_{S}\right]$
    $=P(B=0|F_{S}){\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=0\right)+$
    $\hskip 28.45274ptP(B=1|F_{S}){\operatorname{Var}\,}\left(\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B=1\right)$
    $\leq\left(\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}\right)^{2}+\frac{1}{N^{2+\delta}}\left(\frac{N^{2}}{\rho N}\right)^{2}$ (175)
    where (175) holds by (173), (174), and (156).
  5. 5.
    To show (172), we observe that
    ${\operatorname{Var}\,}\left(\mathbb{E}\left[\frac{(N-\hat{N}_{Y_{ki}})^{2}}{N\hat{N}_{Y_{ki}}}|F_{S},B\right]|F_{S}\right)$
    $\leq P(B=1|F_{S})P(B=0|F_{S})\cdot\left(2\max\left\{\frac{(|\bar{S}|/\rho+\epsilon_{B})^{2}}{N(N_{Q}-\epsilon_{B})}\,,\,\frac{N}{\rho}\right\}\right)^{2}$
    $\leq\frac{1}{N^{2+\delta}}\left(\frac{N}{\rho}\right)^{2}$
    which uses (173), (174), (156), and that the variance of a+(ab)Ua+(a-b)U, where $U\sim\textup{Bernoulli}(p)$ and $a,b$ are scalars, equals $p(1-p)(a-b)^{2}$
This shows (134), (135), and (136), as required by the Lemma.
@@
To prove (137), we observe that for any $i\in[N]$, it holds that
$\sum_{j=1}^{N}1\{|\Gamma_{i}\cup\Gamma_{j}|\neq 0\}$ $=\sum_{j=1}^{N}1\{|\eta_{g(i)}\cup\eta_{g(j)}|\neq 0\}$ (176)
$=\sum_{j=1}^{N}\sum_{m=1}^{M}1\{g(j)=m\}\cdot 1\{|\eta_{g(i)}\cup\eta_{m}|\neq 0\}$ (177)
$=\sum_{m=1}^{M}1\{|\eta_{g(i)}\cup\eta_{m}|\neq 0\}\sum_{j=1}^{N}1\{g(j)=m\}$ (178)
$\leq\sum_{m=1}^{M}1\{|\eta_{g(i)}\cup\eta_{m}|\neq 0\}\cdot\frac{CN}{M}$ (179)
$\leq d^{2}\frac{CN}{M}$ (180)
where (176) holds by (53), (177) and (178) are algebraic manipulations, (179) holds by (50), and (180) holds because
$\sum_{m=1}^{M}1\{|\eta_{g(i)}\cap\eta_{m}|\neq 0\}\leq d^{2}$ (181)
which holds because a group neighborhood $\eta_{g(i)}$ contains at most $d$ groups by (51), each of which are members of at most $d$ group neighborhoods by (52).
To prove (138), we observe that for any $m\in[M]$, it holds that
$\sum_{i=1}^{N}1\{G_{m}\subseteq\Gamma_{i}\}$ $=\sum_{i=1}^{N}1\{m\in\eta_{g(i)}\}$ (182)
$=\sum_{i=1}^{N}\sum_{\ell=1}^{M}1\{g(i)=\ell\}\cdot 1\{m\in\eta_{\ell}\}$ (183)
$=\sum_{\ell=1}^{M}1\{m\in\eta_{\ell}\}\cdot\sum_{i=1}^{N}1\{g(i)=\ell\}$ (184)
$\leq\sum_{\ell=1}^{N}1\{m\in\eta_{\ell}\}\cdot\frac{CN}{M}$ (185)
$\leq d\frac{CN}{M}$ (186)
where (182) holds by (53), (183) and (184) are algebraic manipulations, (185) holds by (50), and (186) holds by (51). ∎
D.5 Proof of Theorem 8
To prove Theorem 8, we will use Lemmas 8-10, and also Theorem 9 which is a local dependence central limit theorem.
Lemma 8 gives a coverage result for $\hat{\tau}_{k}$.
Lemma 8.
Let Assumption 2 hold, and let $\widetilde{V}_{k}(\phi)$ for $\phi\in\{0,1\}^{N}$ and $k=1,2$ be given by (68). It holds for $k=1,2$ that
$P\left(\hat{\tau}_{k}^{\textup{Haj}}-\tau\geq z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{k}}\right)\leq\frac{\alpha}{2}+o(1)$
Lemmas 9 and 10 are used to prove Lemma 8. Lemma 9 gives a central limit theorem for a linearized version of $\hat{\tau}_{k}^{\textup{Haj}}$. Lemma 10 is a corollary of Fatou’s lemma, and will be used to show that $\hat{V}_{k}$ is conservative for the variance of the linearized version of $\hat{\tau}_{k}^{\textup{Haj}}$. It appears as Exercise 3.2.4 in (Durrett, , 2019, 5th edition).
Lemma 9.
Let Assumption 2 hold. For $k\in[2]$, let $\mu_{1}$ and $\mu_{0}$ denote
$\mu_{1}$ $=\sum_{i\in{\mathcal{I}}}\mathbb{E}[\tilde{f}_{ki}(T_{i})],$ $\mu_{0}$ $=\sum_{i\in{\mathcal{I}}}\mathbb{E}[1-\tilde{f}_{ki}(T_{i})].$
where $\tilde{f}_{ki}$ is given by (76). Let $h_{i}$ for $i\in[N]$ denote
where $\phi^{*}$ is given by (102). Let $\sigma^{2}={\operatorname{Var}\,}\sum_{i}h_{i}$, and suppose that $\sigma^{2}\geq N^{2}/M^{4/3-\epsilon}$ for any $\epsilon>0$. Then $\sum_{i}h_{i}/\sigma$ converges to a standard normal.
Lemma 10.
Theorem 9.
(Ross, , 2011, Thm 3.6) Let $U_{1},\ldots,U_{n}$ be random variables such that $\mathbb{E}[U_{i}^{4}]\leq\infty$, $\mathbb{E}U_{i}=0$, with variance denoted by $\sigma^{2}={\operatorname{Var}\,}\sum_{i=1}^{N}U_{i}$, with dependency neighborhoods $N_{1},\ldots,N_{n}$, which are subsets of $[n]$ (which need not be disjoint) such that $U_{j}$ and $U_{i}$ are independent if $U_{j}\notin N_{i}$. Let $D=\max_{i}|N_{i}|$. Let $W=\sum_{i=1}^{N}U_{i}/\sigma$. Then for $Z$ denoting a $N(0,1)$ random variable, and $d_{W}$ denoting Wasserstein metric, it holds that
$d_{W}(W,Z)\leq\frac{D^{2}}{\sigma^{3}}\sum_{i=1}^{n}\mathbb{E}\left[|U_{i}|^{3}\right]+\frac{\sqrt{28}D^{3/2}}{\sqrt{\pi}\cdot\sigma^{2}}\sqrt{\sum_{i=1}^{n}\mathbb{E}\left[U_{i}^{4}\right]}$
Proof of Theorem 8
Proof of Theorem 8.
It holds that
$LB_{1-\alpha}-\tau$ $=\max\left\{\min_{\phi\in\{0,1\}}\hat{\tau}_{1}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}(\phi)}\ ,\ \min_{\phi\in\{0,1\}}\hat{\tau}_{2}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}(\phi)}\right\}-\tau$
$\leq\max\left\{\hat{\tau}_{1}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}}\ ,\ \hat{\tau}_{2}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}}\right\}-\tau$
$=\max\left\{\hat{\tau}_{1}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{1}}-\tau\ ,\ \hat{\tau}_{2}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{2}}-\tau\right\}$ (187)
Let $\mathcal{E}_{k}$ denote the event
$\mathcal{E}_{k}=\left\{\hat{\tau}_{k}-z_{1-\frac{\alpha}{2}}\sqrt{\widetilde{V}_{k}}-\tau\geq 0\right\}.$
$P(\mathcal{E}_{1}\cup\mathcal{E}_{2})\leq\alpha+o(1)$ (188)
Combining (187) and (188) yields that
$P(LB_{1-\alpha}-\tau>0)\leq\alpha+o(1)$ (189)
proving that $LB_{1-\alpha}\leq\tau$ with probability converging to at least $1-\alpha$. ∎
Proof of Lemmas 8 and 9
Proof of Lemma 8.
Let $h_{i}$ be defined as in Lemma 9, and let $\sigma^{2}={\operatorname{Var}\,}\sum_{i}h_{i}$. Divide the sequence of experiments into two subsequences, depending on whether $\sigma^{2}\geq N^{2}/M^{4/3-\epsilon}$.
For the subsequence that satisfies $\sigma^{2}\geq N^{2}/M^{4/3-\epsilon}$, it holds by linearization arguments made in the proof of Theorem 6 that
$\frac{\hat{\tau}_{k}^{{\textup{Haj}}}-\tau}{\sigma}$ $=\frac{\nabla h(\mathbb{E}\xi)^{T}(\xi-\mathbb{E}\xi)}{\sigma}+O_{P}\left(\frac{1}{\sigma}\right)$ (190)
@@ D.5 Proof of Theorem 8
$\frac{\hat{\tau}_{k}^{{\textup{Haj}}}-\tau}{\sigma}\rightarrow N(0,1)$
@@
$\liminf\limits_{N\rightarrow\infty}{\operatorname{Var}\,}\left[\frac{\hat{\tau}_{k}^{\textup{Haj}}-\tau}{\sigma}\right]\geq 1$ (192)
@@
$\frac{\hat{V}_{k}}{\sigma^{2}}$ $=\frac{\mathbb{E}\hat{V}_{k}}{\sigma^{2}}+\frac{1}{\sigma^{2}}O_{P}\left(\frac{N^{2}\log M}{M^{3/2}}\right)$
$\geq{\operatorname{Var}\,}\left(\frac{\hat{\tau}_{k}^{\textup{Haj}}}{\sigma}\right)+o_{P}\left(1\right)$ (193)
@@
$\liminf\limits_{N\rightarrow\infty}\frac{\hat{V}_{k}}{\sigma^{2}}\geq 1$
@@
$P\left(\hat{\tau}_{k}^{\textup{Haj}}-\tau\geq z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{k}}\right)\leq\frac{\alpha}{2}+o(1)$ (194)
for the subsequence of experiments that satisfies $\sigma^{2}\geq N^{2}/M^{4/3-\epsilon}$.
@@
$P\left(\left|\sum_{i}h_{i}\right|\geq\frac{N}{M^{2/3-\epsilon}\sqrt{\alpha/2}}\right)\leq\alpha/2$
where we have used that $\frac{N}{M^{2/3-\epsilon}}=\omega(\sigma)$ for $\epsilon>0$. In turn, this implies that
$P\left(\hat{\tau}_{k}-\tau\geq\frac{N}{M^{2/3-\epsilon}\sqrt{\alpha/2}}\right)\leq 1-\alpha+o(1)$ (195)
where we have used that $\hat{\tau}_{k}-\tau=\sum_{i}h_{i}+O_{P}(1)$, which holds by (190). Applying (195) yields
$P\left(\hat{\tau}_{k}-\tau\geq z_{1-\frac{\alpha}{2}}\sqrt{\Delta}\right)\leq\alpha/2+o(1)$ (196)
where $\Delta=N/(M^{2/3-\epsilon}z_{1-\frac{\alpha}{2}}^{2}\cdot\alpha/2)$. Combining (194) and (196) yields that for all experiments, it holds that
$P\left(\hat{\tau}_{k}-\tau\geq z_{1-\frac{\alpha}{2}}\sqrt{\max\left(\hat{V}_{k},\Delta\right)}\right)\leq\alpha/2+o(1),$
As $\widetilde{V}_{k}$ can be seen to equal $\max\left(\hat{V}_{k},\Delta\right)$, this proves the lemma.
@@
Let $U_{m}$ for $m\in[M]$ be given by
$U_{m}=\sum_{i\in G_{m}}h_{i}$
Since $|h_{i}|\leq\frac{1}{\rho}$ and $|G_{m}|\leq CN/M$, it follows that $|U_{m}|\leq\frac{CN}{M\rho}$. Additionally, it can be seen that the dependency neighborhoods of $U_{1},\ldots,U_{M}$ have maximum size $d^{2}$, since each $U_{m}$ depends on $X$ only through $\{X_{G_{\ell}}:\ell\in\eta_{m}\}$. As a result, applying Theorem 9 yields
$d_{W}\left(\frac{\sum_{i}h_{i}}{\sigma}\ ,\ N(0,1)\right)$ $\leq\frac{d^{4}}{\sigma^{3}}\cdot M\left(\frac{CN}{M\rho}\right)^{3}+\frac{\sqrt{28}}{\sqrt{\pi}}\cdot\frac{d^{3}}{\sigma^{2}}\sqrt{M\left(\frac{CN}{M\rho}\right)^{4}}$
$=O\left(\frac{N^{3}}{\sigma^{3}M^{2}}\right)+O\left(\frac{N^{2}}{M^{3/2}\sigma^{2}}\right)$ (197)
@@
where (198) uses the assumption of the lemma that $\sigma^{2}\geq\frac{N}{M^{2/3-\epsilon}}$ for $\epsilon>0$. ∎